Need help with question 1
Write the balanced equation for the reaction betwe
Answer #1

Hi, The equation you have is correct. The mole ratio of acid to base is 1:2 since 1 mol of acid reacts with 2 moles of base.

for c, Molarity is given by: M = \frac{\text{moles}}{\text{Volume in L}} so the number of moles of base are given by:

\text{moles of base} = M_{b} \times V_{b}

the ratio of acid to base is \frac{a}{b} so the moles of acid are:

\text{moles of acid} = \frac{a\times M_{b} \times V_{b}}{b} I hope you get the idea of where the equation comes from. Now what you have to do it to put the values: a=1, b=2. Mb is the molarity of sodium hydroxide that you have used in the experiment and Vb is the volume that you have transferred in the flask to complete the titration. For example you used 0.1 M NaOH and the volume consumed was 13.2 mL the number of moles of acid would be:

\text{moles of acid} = \frac{1\times 0.1 M \times 0.0132 L}{2} = 6.6\times10^{-4} mol

You need to put your own values from the lab experiment.

Now, you have to divide the moles of acid by the mass of candy

suppose you had a 7 grams candy so

\frac{6.6\times10^{-4} mol}{7 g} = 9.43 \times 10^{-5} \text{mol acid/g candy}

Hope that helps.